Shift summary for OWL shift 9/Jan/2000


Subject: Shift summary for OWL shift 9/Jan/2000
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:11:26 CST


                     Shift summary for owl shift 9/Jan/2000

On Shift: NS, AG

Hours of beam available at all intensities: 6.8
Hours of beam available at nominal intensity: 6.8
Hours of E799 data (special runs not included): 5.8
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:

o Rebooting pipe-line crates, one dart_fresh_start.

o Resetting pipe-line + kfrends.

o Minor accelerator glitch.

o Changing tapes.

========================================================================

Tapes brought to FCC this shift:

UP{A,B,C}31[0-2], UPF07[4-5], UPT025, UPM055, UPS104, UPH055

========================================================================

Accesses during shift: none.

========================================================================

Runs ended this shift:

15443 - 7.8M - BEAM_799
15444 - 13.2M - BEAM_799
 
=======================================================================

Overview of shift:

Very smooth E799 data-taking after pipe-line crates were rebooted at
the beginning of the shift.

Mounted tapes UPE021, UPK040, UPH054 as per ONSPLIT request.

========================================================================

Description of Shift:

run spill time comment
------------------------------------------------------------------------

15443 0000 Run is ongoing as shift starts. "Message from the
                        pipe-line" shows up every now and then.
                 
                0005 Pipe-line errors, resetting doesn't work - hangs
                        after crate 8. Rebooting crates generates errors.
                        Stop run 15443. Clean pipe. Do dart_fresh_start
                        -> all pipe-line crates come up red upon
                           initialization.

                0035 Try rebooting pipe-line crates one at a time...
                        now it works -> reboot all crates, initialize
15444 0053 pipe-line, start run 15444.

15444 108 0318 Beam is gone. Linac Klystron buncher tripped; ops
                        investigating.

         117 0326 Beam is back.

         260 0640 Channels 960 and 964 are consistently not
                        illuminated for a few spills in a raw. Also
                        getting "Ped problem ..." from Val.
                        -> Reset pipe-line w/DBS, lost one spill, but that
                           fixes it.

                0652 Tapes are full, although KTeV DAQ monitoring shows
                       barely more than 50% (?) ... Run stops.
                       -> Change tapes. Initialize new run.
                
15445 0703 Run 15445 started.

========================================================================
Items to Follow up on Next Shift

  Keep an eye on DCOR LV power supply for DC2, it is reading low (4.8V).

  Label 30 QKE tapes for short reg run and other special runs

  Don't forget to hit "reset quartz" button after each KFREND.



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:11:28 CST